### v1

@@ 2 Nuclear EOS models
In this model, matter in a Wigner–Seitz (WS) cell is composed of free nucleons, alpha particles, and a representative heavy nucleus. The nucleon number density averaged over the WS cell is defined by
$n_{\mathrm{B}}=\frac{1}{V_{\mathrm{cell}}}\int_{\mathrm{cell}}d\mbox{\boldmath$r$}\,\left[n_{\mathrm{p}}(r)+n_{\mathrm{n}}(r)+4n_{\alpha}(r)\right],$ (1)
where $V_{\mathrm{cell}}$ is the volume of the WS cell, and $n_{i}(r)$ denotes the local number density of species $i$ at position $\boldsymbol{r}$ inside the cell. The average proton fraction within the cell is defined by
$Y_{\mathrm{p}}=\frac{1}{n_{\mathrm{B}}V_{\mathrm{cell}}}\int_{\mathrm{cell}}d\mbox{\boldmath$r$}\,\left[n_{\mathrm{p}}(r)+2n_{\alpha}(r)\right].$ (2)
@@
The nuclear saturation parameters used in this study are summarized in Table 1. These EOS models are taken from Ref. 2023PTEP.2023f3D03O , and correspond to different choices of the symmetry energy slope parameter $L$, while keeping the incompressibility $K$ fixed. The saturation density $n_{0}$, the saturation energy $w_{0}$, and the symmetry energy $S_{0}$ are optimal values. The labels of the EOS models, except for EOS Z, follow those used in Ref. 2007PhRvC..75a5801O .
@@
for $n_{\rm B}>n_{0}$. In the present study, each EOS listed in Table 1 is extended to supranuclear densities using common saturation parameters, such as $S_{0}$ and $L$. Meanwhile, the parameter $S_{00}$, newly introduced in Eq. (14), represents the symmetry energy at a density of $2n_{0}$, which is typically assumed to be around 45 MeV 2019EPJA…55…39Z ; 2020PhRvR…2b2033L ; 2020PhRvL.125t2702D . In addition, EOS models with higher $S_{0}$ and larger $L$ tend to predict higher symmetry energy at supranuclear densities. Thus, as for the reference model, we choose $S_{00}=60$ MeV for EOS B ($L=73.4$ MeV), $S_{00}=50$ MeV for EOS E ($L=42.6$ MeV), $S_{00}=40$ MeV for EOS Z ($L=23.7$ MeV), and $S_{00}=30$ MeV for EOS H ($L=7.6$ MeV). So as to assess the effect of $S_{00}$, we also examine models with $S_{00}$ ranging from 40 to 60 MeV for EOS E and from 35 to 50 MeV for EOS Z. In the following, models are denoted as Xyy, where the letter X identifies the subnuclear-density EOS (B, E, Z or H), and yy indicates the assigned value of $S_{00}$. Furthermore, we extend the EOS at supranuclear densities to finite temperatures in a manner consistent with the subnuclear regime and write the free energy per baryon as

### v2

@@ 2 Nuclear EOS models
In this model, matter in a Wigner–Seitz (WS) cell is composed of free nucleons, alpha particles, and a representative heavy nucleus. The local number densities of protons and neutrons, $n_{\mathrm{p}}(r)$ and $n_{\mathrm{n}}(r)$, at position $\boldsymbol{r}$ inside the cell include contributions from both free nucleons and nucleons bound inside the heavy nucleus. Alpha particles are treated as a separate species, with local number density $n_{\alpha}(r)$.
With these definitions, the nucleon number density averaged over the WS cell is defined by
$n_{\mathrm{B}}=\frac{1}{V_{\mathrm{cell}}}\int_{\mathrm{cell}}d\mbox{$r$}\,\left[n_{\mathrm{p}}(r)+n_{\mathrm{n}}(r)+4n_{\alpha}(r)\right],$ (1)
where $V_{\mathrm{cell}}$ is the volume of the WS cell. Thus, although the presence of the heavy nucleus is not explicitly visible in Eq. (1), the integration over the WS cell properly accounts for all nucleons in the system. The average proton fraction within the cell is defined by
$Y_{\mathrm{p}}=\frac{1}{n_{\mathrm{B}}V_{\mathrm{cell}}}\int_{\mathrm{cell}}d\mbox{$r$}\,\left[n_{\mathrm{p}}(r)+2n_{\alpha}(r)\right].$ (2)
@@
The nuclear saturation parameters used in this study are summarized in Table 1. These EOS models are taken from Ref. 2023PTEP.2023f3D03O , and correspond to different choices of the symmetry energy slope parameter $L$, while keeping the incompressibility $K$ fixed at 230 MeV. The saturation density $n_{0}$, the saturation energy $w_{0}$, and the symmetry energy $S_{0}$ are optimal values. The labels of the EOS models, except for EOS Z, follow those used in Ref. 2007PhRvC..75a5801O .
@@
for $n_{\rm B}>n_{0}$. In the present study, each EOS listed in Table 1 is extended to supranuclear densities using common saturation parameters, such as $S_{0}$ and $L$. Meanwhile, the parameter $S_{00}$, newly introduced in Eq. (14), represents the symmetry energy at a density of $2n_{0}$, which is typically assumed to be around 45 MeV 2019EPJA…55…39Z ; 2020PhRvR…2b2033L ; 2020PhRvL.125t2702D . In addition, EOS models with higher $S_{0}$ and larger $L$ tend to predict higher symmetry energy at supranuclear densities. Thus, as for the reference model, we choose $S_{00}=60$ MeV for EOS B ($L=73.4$ MeV), $S_{00}=50$ MeV for EOS E ($L=42.6$ MeV), $S_{00}=40$ MeV for EOS Z ($L=23.7$ MeV), and $S_{00}=30$ MeV for EOS H ($L=7.6$ MeV). So as to assess the effect of $S_{00}$, we also examine models with $S_{00}$ ranging from 40 to 60 MeV for EOS E and from 35 to 50 MeV for EOS Z. In the following, models are denoted as Xyy, where the letter X identifies the subnuclear-density EOS (B, E, Z or H), and yy indicates the assigned value of $S_{00}$. Incidentally, as already mentioned, the incompressibility is fixed at $K=230$ MeV for all EOS models. Furthermore, we extend the EOS at supranuclear densities to finite temperatures in a manner consistent with the subnuclear regime and write the free energy per baryon as